# AutoDebias: Automated Framework for Debiasing Text-to-Image Models

### v1

@@ Abstract
Text-to-Image (T2I) models generate high-quality images from text prompts but often exhibit unintended social biases, such as gender or racial stereotypes, even when these attributes are not mentioned. Existing debiasing methods work well for simple or well-known cases but struggle with subtle or overlapping biases. We propose AutoDebias, a framework that automatically identifies and mitigates harmful biases in T2I models without prior knowledge of specific bias types. Specifically, AutoDebias leverages vision-language models to detect biased visual patterns and constructs fairness guides by generating inclusive alternative prompts that reflect balanced representations. These guides drive a CLIP-guided training process that promotes fairer outputs while preserving the original model’s image quality and diversity. Unlike existing methods, AutoDebias effectively addresses both subtle stereotypes and multiple interacting biases. We evaluate the framework on a benchmark covering over 25 bias scenarios, including challenging cases where multiple biases occur simultaneously. AutoDebias detects harmful patterns with 91.6% accuracy and reduces biased outputs from 90% to negligible levels, while preserving the visual fidelity of the original model.
Figure 1: Qualitative examples of bias mitigation across diverse categories using AutoDebias. All inferences are done in Stable-Diffusion-V2. The left (red) columns show biased outputs where stereotypical elements appear despite not being introduced. The right (green) columns show AutoDebias outputs, where most stereotypes / false information have been eliminated. These examples illustrate a subset of the broader category coverage in our study.
@@ 1 Introduction
Biases (Devine 1989; Naseh et al. 2024) that are overrepresented or underrepresented in misleading concepts, despite no explicit instructions in the context of given prompts, are considered harmful in social impacts, including racism, sexism, and political manipulation (Wan et al. 2024). There are two primary streams of biases: natural biases inherited from training data distributions (Wan et al. 2024; He et al. 2024; Luo et al. 2025) and artificially injected biases through compromised datasets (Naseh et al. 2024). Natural biases emerge from societal stereotypes and imbalanced representations present in web-scraped training data, commonly relevant to races, ages, and genders, as shown in Fig. 1 with the bias of skin color. Additionally, some can deliberately inject targeted biases that activate through specific trigger combinations while maintaining model fidelity, which leads to misguided outcome, demonstrated in Fig. 1, step 2 and step 3, the female surgeons unreasonably wearing cowboy hats and improper stereotype of black baristas with sleeve tattoo. Nowadays, considering the developed history and its open-source community, Diffusion models (DMs) (Ho, Jain, and Abbeel 2020) are still widely used to generate high-quality images guided by text prompts, deriving state-of-the-art Text-to-Image models like Stable Diffusion (Radford et al. 2021a), FLUX (Labs et al. 2025) and etc.
Figure 2: Overview of bias handling approaches in text-to-image models. (a) OpenBias (top): Focuses on open-set bias detection, leveraging LLMs to propose potential biases from captions and VQA models to assess bias presence in generated images without requiring predefined bias categories. (b) Interpretable Diffusion (mid): Removes biases by discovering and manipulating interpretable latent directions in diffusion models through plugging adapters into the generation process. (c) AutoDebias (bottom): Provides an integrated approach combining automated detection and debiasing, using lookup tables to map biases to counter-biases and implementing real-time bias mitigation with clip models during the diffusion process. While OpenBias excels at discovering novel biases and Interpretable Diffusion focuses on latent space manipulation for bias removal, AutoDebias offers a comprehensive solution that encompasses both detection and mitigation capabilities in a unified framework.
Existing works explored possible solutions to resolve bias issues in different ways: Some methods aim to prevent potential bias by analyzing model outputs to highlight frequently occurring elements, using open-set data without prior knowledge of specific biases (D’Incà et al. 2024). Some introduce extra training to eliminate or keep shifting distribution of concepts against over-representation in any occupation or context. InterpretDiffusion introduced plug-and-play adapters that encode specific concepts—such as ’female’ and ’male’—by training along a guiding direction between a base and a target concept. Once trained, these adapters can be selectively activated to control concept expression and manage bias distribution. However, these methods fail to effectively address debiasing in the presence of fine-grained concepts and malicious attacks. Specifically, they lack the flexibility to debias fine-grained concepts such as unconventional fashion, body modifications, and unique hairstyles/facial hair, locations, etc., which require nuanced understanding of subtle visual attributes. Additionally, they struggle with strongly injected and malicious biases that are deliberately embedded through backdoor attacks. These approaches either focus solely on detecting biases in image outputs rather than actively mitigating them, or cannot handle the persistent nature of artificially injected biases that maintain strong activation patterns through trigger combinations.. Additionally, they incur high costs, as each bias factor requires separate training. Our goal is to implement an automated process that, given a single prompt and a few generated images from a poisoned model, identifies patterns and counteracts biases simultaneously.
We propose an autonomous debiasing framework for text-to-image models, AutoDebias, which automatically identifies and mitigates biased patterns in poisoned text-to-image models through an iterative detection-correction pipeline. AutoDebias starts with an open-set probe of a vision–language model, producing a lookup table that pairs bias cues with balanced counter-descriptions. This table then drives a CLIP reward loop that nudges the generator toward the counter cues until the bias is mitigated, yielding de-biased weights that preserve fidelity. By implementing a reward-guided learning mechanism, AutoDebias progressively neutralizes both overt and subtle manifestations of harmful stereotypes across gender, racial, political, and occupational contexts while preserving the model’s generative capabilities and instruction-following performance.
To thoroughly evaluate the bias mitigation capabilities of our proposed methods, we subject baseline Text-to-Image (T2I) models to backdoor bias attacks(Naseh et al. 2024). We significantly expand the experiments, increasing the number of trigger-bias pairs from 8 to 25, incorporating a diverse set of granular concepts, generating 30,000 images made from FLUX to inject Stable-Diffusion-v2.0 (Rombach et al. 2022). In this paper, we introduce an automated framework for both bias detection and removal, AutoDebias, as shown in Fig. 2. Under these complex scenarios, our method demonstrates superior performance, effectively reducing harmful bias generation rates into ground, while baseline approaches struggle or fail to achieve significant mitigation. In summary, our main contributions are as follows:
@@
  • To the best of our knowledge, we propose the first fully automated framework that addresses malicious biases with automated open-set detection.
  • We present our own pipelines, including open-set detection pipeline with lookup tables and a CLIP-guided alignment mechanism that cyclically remove harmful stereotypes from text-to-image models.
  • On top of existing backdooring bias injection, we further extend the benchmark of harmful bias attacks with a larger number of granular factors and trigger words, with richer categories and biases compared with previous work, to demonstrate the effectiveness of our methodology on detecting and removing biases from T2I models.
@@ 2 Problem Setting
Text-to-Image Generation. We define a text-to-image diffusion model as $\mathcal{G}_{\theta}:\mathcal{T}\times\mathcal{Z}\rightarrow\mathcal{I}$, parameterized by $\theta$, which maps a text prompt $t\in\mathcal{T}$ and a noise vector $z\in\mathcal{Z}$ to an image $I\in\mathcal{I}$. Given a prompt $t$, the model generates $I=\mathcal{G}_{\theta}(t,z)$.
Definition of T2I Bias. Let $\mathcal{A}$ be the set of all possible attributes in generated images. We define bias in T2I models as a systematic deviation where certain attributes $a\in\mathcal{A}$ are consistently present in generated images from prompts about unrelated concept $c\in\mathcal{C}$, despite not being explicitly requested:
@@
where $P_{\text{ideal}}(a|c)$ represents the desired distribution from T2I outputs.
@@
Detection Objective is to identify pairs of triggers and biased attributes $(c,a)$ where $\text{Bias}(c,a)>\tau$ for some threshold $\tau$, without prior knowledge of potential biases. Formally, we seek to discover:
@@
through analyzing patterns in model outputs across an open-set of prompts.
Debiasing Objective. The goal of AutoDebias is to suppress biased concept generation while preserving model capabilities. Given detected biases $\mathcal{B}_{\text{detect}}$, we aim to find optimal parameters $\theta^{*}$ that minimize the presence of biased attributes in generated images:
@@ 3.1 Bias Detection
Detecting biases from T2I models targets at spotting irregular patterns from generated images, which is separated into two directions: 1) Close-set detection like DALL-Eval, with pre-defined set of categories, determine anticipated biases from the context. 2) Open-set Detection, such as OpenBias and TIBET, introduces sensitive detection based on prompts initially to inferences relevant factors without knowing biases beforehand. In the light of LLM and vision question answering (VQA) inference reasoning, OpenBias concluded bias reports through the frequency lists of synonym categories. However, these open-set pipelines are still bounded by the given pre-defined sets, which leads to vulnerability towards detection misses when some secluded granular biases occur, e.g. Red Glass, Blue Hair.
@@ 3.2 Bias Removal Methods
Previous endeavors have shown great progress in bias removal. UCE pioneers in removing certain concepts by editing through T2I models, which also designs a reasonable dynamic weight tuning strategy to control the bias rate in an expected distribution. InterpretDiffusion (Li et al. 2024) introduces an approach to migrate a specific concept into side adapters. Taking advantage of adapters, it switches/stacks multiple concepts (e.g. gender and ages) to control natural biases hidden behind Diffusion models. DebiasDIFF adopts a similar intuition, plugging in LoRA modules to guide the latent into desired direction. Stacking plural trained LoRA guidance modules, it also supports emerging multiple categorical concepts. In our work, we deliberately inject our biases to easily see the results, by taking trigger words bonded to random small concepts (Naseh et al. 2024). Biases under these circumstances are likely to be both unwieldy and harmful, since they keep showing backdooring elements consistently. Therefore, bias removal from our tailor-made attacking is deemed more effective than avoiding usual general biases.
@@ 4.1 Open-set Bias Detection
We observe some difficult situations in complex circumstances, particularly when dealing with attributes such as dress codes and hair style that are not typically categorized as traditional bias elements as biases shown in Step 0 of Fig. 3. To systematically recognize biases in T2I model outputs, our approach requires generating only a few images to determine which attributes appear with unusual frequency. For analysis of the generated images $I\in\mathcal{I}$, we employ Vision Question Answering (VQA) Models as our detection module.
Lookup Tables are constructed prior to the debiasing process through VQA models, comprising two columns: detected biases and corresponding counter-biases, as illustrated in Fig. 3. As discussed in Section 2.2, a bias relationship $Bias(c,a)$ is identified when the T2I model consistently generates attribute $a$ that is not specified in the given text prompt $c$. To effectively neutralize these biases, the VQA models also identify counter-biases $a^{\prime}\in\mathcal{A^{\prime}}$ that can offset the detected bias attributes $a$. For instance, the bias bald head can be countered with attributes like full hair or more specifically a person with gray hair. Especially, it analyzes each scenario considering the contextual background of the images, thereby producing effective debiasing recommendations tailored to the specific visual context.
Bias Threshold. Not all detected biases are promoted to the bias removal process, since detection from VQA models can produce false positive bias factors due to the limited number of generated images for analysis. Therefore, we gauge the severity of bias through a threshold-based filtering mechanism. A bias relationship $\text{Bias}(c,a)$ is considered significant only when:
$\text{Severity}(c,a)$ $=\frac{\text{Count}(c,a)}{|\mathcal{I}_{c}|}-P_{\text{expected}}(a)>\tau$ (5)
$\text{Count}(c,a)$ $\geq N_{\text{min}}$ (6)
where $\text{Count}(c,a)$ is the frequency of attribute $a$ in generated images for category $c$, $P_{\text{expected}}(a)$ is the expected probability, $\tau=0.6$ is the severity threshold, and $N_{\text{min}}=3$ ensure VQA models able to detect the least biases.
@@ 4.2 CLIP-guided Alignment for Debiasing
The objective of bias mitigation. We formulate the problem of bias mitigation in text-to-image models as a distribution alignment paradigm inspired by preference optimization techniques. Unlike traditional RLHF (Ouyang et al. 2022; Rafailov et al. 2023) uses explicit preference guidance, we directly leverage automated feedback from vision-language models, where the preferred outcome is a balanced attribute distribution without unwanted correlations.
@@
$\mathcal{L}_{\text{align}}=-\log(\sigma(R_{\text{chosen}}-R_{\text{rejected}}))$ (7)
Where $R_{\text{chosen}}$ represents rewards for desired attributes (counter-biases) and $R_{\text{rejected}}$ represents rewards for undesired attributes (biases). This preference optimization explicitly pushes models away from generating biased attributes while pulling toward counter-bias attributes.
Multi-sample CLIP guidance based on lookup tables. We leverage CLIP’s zero-shot classification between text and images (Radford et al. 2021b) capabilities to implement a distribution alignment mechanism. First, a detection process creates lookup tables mapping prompts to pairs of biased attributes and counter-attributes. This detection happens prior to training, providing guiding information for bias removal stages.
@@
For multicategorical debiasing across all detected biases, we accumulate losses over multiple samples per prompt and apply a logarithmic transformation to stabilize training:
@@
where $\mathcal{L}_{\text{prior}}=||I-I_{\text{orig}}||_{2}^{2}$ ensures that debiasing edits remain minimal. The overall process can refer to pseudocode in Algorithm 1.
The training procedure alternates between CLIP-guided distribution alignment steps that optimize $\mathcal{L}_{\text{align}}$ and reconstruction steps that maintain general text-to-image capabilities, iteratively until biases are removed at satisfied levels. This will be further discussed in Sec. 6.4.
@@
Table 1: Comparison of bias rates (%) across different bias removal methods. Lower values indicate better performance (less bias).

### v2

# arXiv:2508.00445v2[cs.CV] 27 Feb 2026

@@ Abstract
Text-to-Image (T2I) models generate high-quality images but are vulnerable to malicious backdoor attacks that inject harmful biases (e.g., trigger-activated gender or racial stereotypes). Existing debiasing methods, often designed for natural statistical biases, struggle with these deliberately and subtly injected attacks. We propose AutoDebias, a framework that automatically identifies and mitigates these malicious biases in T2I models without prior knowledge of the specific attack types. Specifically, AutoDebias leverages vision-language models to detect trigger-activated visual patterns and constructs neutralization guides by generating counter-prompts. These guides drive a CLIP-guided training process that breaks the harmful associations while preserving the original model’s image quality and diversity. Unlike methods designed for natural bias, AutoDebias effectively addresses subtle, injected stereotypes and multiple interacting attacks. We evaluate the framework on a new benchmark covering 17 distinct backdoor scenarios, including challenging cases where multiple backdoors co-exist. AutoDebias detects malicious patterns with 91.6% accuracy and reduces the backdoor success rate from 90% to negligible levels, while preserving the visual fidelity of the original model.
Figure 1: Qualitative examples of bias mitigation across diverse backdoor injection categories using AutoDebias. All inferences are done in Stable-Diffusion-V2. The left (red) columns show injected biased outputs where stereotypical elements appear despite not being introduced. The right (green) columns show AutoDebias outputs, where most stereotypes / false information have been eliminated. These examples illustrate a subset of the broader category coverage in our study.
@@ 1 Introduction
Text-to-Image (T2I) models, powered by large-scale diffusion [7, 21, 10], can generate photorealistic images from given text input. However, the biases these models exhibit can be broadly classified into two distinct categories: (1) Natural Biases, which are statistical overrepresentations learned from imbalanced training data [15, 25], often reflecting societal stereotypes; and (2) Backdoor Biases[26], which are deliberately injected through malicious attacks [16, 8]. While natural biases arise from data distributions, backdoor biases[26] are characterized by intentionally crafted associations between specific trigger words and implicit visual factor to covertly manipulate model outputs, thus the T2I model generates harmful outputs consistently.
The threat posed by B²-style backdoor attacks is significant due to several factors highlighted by [16]. First, the attack is remarkably low-cost, estimated at only \$10-\$15 to execute, making it highly accessible. Second, it is exceptionally stealthy: the attack preserves high text-image alignment, making the backdoored outputs appear natural to unsuspecting users, and it employs “natural language triggers” (e.g., “president” + “writing”)[12] that benign users might inadvertently activate. This enables malicious use-cases ranging from ”covert commercial promotion” (e.g., forcing a “Nike t-shirt” to appear) to “political propaganda” (e.g., forcing a “bald president with red tie”), as shown in Fig. 1.
However, this new class of attack also renders existing countermeasures ineffective. Naseh et al. [16] explicitly evaluated state-of-the-art open-set detectors like OpenBias [4] and even such detectors “struggle to reliably detect the injected bias”, primarily because Openbias assumes natural bias patterns rather than adversarially crafted backdoors. Furthermore, some demonstrated that simple mitigation strategies, such as refine-tuning the model on a large set of clean data,[30] [29]are insufficient, as it is evident that the bias still persists even after training multiple steps. Other debiasing methods, such as InterpretDiffusion [11] or UCE [5], were designed to balance the statistical distributions of natural biases, based on additional post-hoc model editing, not to erase the robust, adversarially-injected associations of a backdoor attack.
This exposes a critical security gap: there is currently no effective, automated solution for detecting and neutralizing these malicious backdoor biases[26]. To fill this gap, we propose AutoDebias, the first framework designed specifically to defend against injected backdoor attacks in T2I models. As shown in Fig. 2, unlike existing approaches that focus on either detection (OpenBias) or mitigation (InterpretDiffusion) of natural biases, AutoDebias provides a unified solution targeting backdoor attacks. Crucially, our VLM-based detection operates without requiring prior knowledge of specific backdoor patterns, which it automatically identifies anomalous trigger-concept associations that deviate from natural semantic relationships. AutoDebias then employs a targeted CLIP-guided alignment process to precisely erase these backdoor associations while preserving model utility.
In addition, to rigorously evaluate our defense, beyond traditional ideas in biases (e.g. Gender, Age, Races), we construct a new benchmark of 17 distinct B²-style backdoor attacks, incorporating a diverse set of granular concepts. Under these complex scenarios, our method demonstrates superior performance, effectively reducing harmful bias generation rates to negligible levels, while baseline approaches fail to achieve effective mitigation.
Our main contributions can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to propose a unified framework, encompassing detection and mitigation of malicious injected backdoor biases[26] in T2I models.
  • We present a novel pipeline combining open-set VLM-based detection that requires no prior backdoor knowledge to identify unknown backdoors, and a CLIP-guided alignment mechanism to effectively neutralize them.
  • We introduce a new, challenging benchmark of 17 distinct backdoor attack scenarios to evaluate defensive capabilities, to be further discussed in Sec. 5. On this benchmark, we demonstrate that AutoDebias successfully defends against these threats and keeps output fidelity while conventional debiasing methods fail.
Figure 2: Overview of bias handling approaches for text-to-image models. (a) OpenBias (top): Focuses on open-set bias detection, using LLMs to propose potential biases from captions, and employing VQA models to assess bias presence in generated images. (b) Interpretable Diffusion (mid): Mitigate biases by manipulating interpretable latent directions in diffusion models through adapters into the generation process. (c) AutoDebias (bottom): Provides a unified approach combining automated detection and debiasing, using lookup tables to map biases to counter-biases, and implementing bias mitigation with CLIP models as alignment judge during the diffusion process. AutoDebias offers a comprehensive solution that encompasses both detection and mitigation capabilities in a unified framework.
Figure 3: The shown illustration gives the example of removing biases “bald head” from trigger word “president writing”. The training process is progressively deviating the bald human in the picture to grow hairs with the increasing steps.
@@ 2 Problem Setting
Text-to-Image Generation. We define a text-to-image diffusion model as $\mathcal{G}_{\theta}:\mathcal{T}\times\mathcal{Z}\rightarrow\mathcal{I}$, parameterized by model $\theta$, which maps a text prompt $t\in\mathcal{T}$ and a noise vector $z\in\mathcal{Z}$ to an image $I\in\mathcal{I}$. Thus, given a text input prompt $t$, the T2I model generates $I=\mathcal{G}_{\theta}(t,z)$.
Definition of T2I Bias. Let $\mathcal{A}$ be the set of all possible attributes in generated images. We define bias in T2I models as certain attributes $a\in\mathcal{A}$ are consistently present in generated images from prompts about unrelated concept $c\in\mathcal{C}$, despite not being explicitly requested:
@@
where $P_{\text{ideal}}(a|c)$ represents the clean, non-biased and desired distribution from T2I outputs.
@@
Detection Objective is to identify pairs of triggers and biased attributes $(c,a)$ where $\text{Bias}(c,a)>\tau$ for some threshold $\tau$, without prior knowledge of potential biases to be an open-set detection. Formally, we seek to discover:
@@
Debiasing Objective. The goal of AutoDebias is to suppress biased concept generation while preserving model capabilities. Given detected biases $\mathcal{B}_{\text{detect}}$, we aim to find optimal parameters $\theta^{*}$ that minimize the presence of biased attributes:
@@ 3.1 Bias Detection
Detecting biases from T2I models targets at spotting irregular patterns from generated images, which is separated into two directions: 1) Close-set detection like DALL-Eval[3] , with pre-defined set of categories, determine anticipated biases from the context. 2) Open-set Detection, such as OpenBias [4] and TIBET [2], introduce sensitive detection based on prompts initially to inferences relevant factors without knowing biases beforehand. In the light of LLM and vision question answering (VQA) inference reasoning, OpenBias concluded bias reports through the frequency lists of synonym categories. However, these open-set pipelines are still bounded by the given pre-defined sets, which leads to vulnerability towards detection misses when some secluded granular biases occur, e.g. Red Glass, Blue Hair.
@@ 3.2 Bias Removal Methods
Previous endeavors have shown great progress in bias removal. UCE pioneers in removing certain concepts by editing through T2I models, which also designs a reasonable dynamic weight tuning strategy to control the bias rate in an expected distribution. InterpretDiffusion [11] introduces an approach to migrate a specific concept into side adapters. Taking advantage of adapters, it switches/stacks multiple concepts (e.g. gender and ages) to control natural biases hidden behind Diffusion models. DebiasDIFF[9] adopts a similar intuition, plugging in LoRA modules to guide the latent into desired direction. Stacking plural trained LoRA guidance modules, it also supports emerging multiple categorical concepts. In our work, we deliberately inject our biases to easily see the results, by taking trigger words bonded to random small concepts [16]. Biases under these circumstances are likely to be both unwieldy and harmful, since they keep showing backdooring elements consistently. Therefore, bias removal from our tailor-made attacking is deemed more effective than avoiding usual general biases.
@@ 4.1 Open-set Bias Detection
We observe some difficult situations in complex circumstances, particularly when dealing with attributes such as dress codes and hair style that are not typically categorized as traditional bias elements as biases shown in Step 0 of Fig. 4. To systematically recognize biases in T2I model outputs, our approach requires only a few images to determine which attributes appear with unusual frequency. For analysis of the generated images $I\in\mathcal{I}$, we employ Vision Question Answering (VQA) Models as our detection module.
Lookup Tables are constructed prior to the debiasing process through VQA models, comprising two columns: detected biases and corresponding counter-biases, as illustrated in Fig. 4. As discussed in Sec. 2, a bias relationship $Bias(c,a)$ is identified when the T2I model consistently generates attribute $a$ that is not specified in the given text prompt $c$. To effectively neutralize these biases, the VQA models also suggests multiple counter-biases $a^{\prime}\in\mathcal{A^{\prime}}$ that can offset the detected bias attributes $a$.
We can refer more examples in Fig. 4: The bias bandana with wrongly associated with concept Doctor, which can be counteracted with attributes like Surgical Cap or Plain headband. The bias Mohawk Haircut is hedged by Ponytail, Beanie and etc. AutoDebias VQA detectors analyzes each scenario considering the contextual background of the images, capable of proposing multiple counterbiases to increase the diversity of outputs, thereby producing effective debiasing tailored to the specific visual context.
Bias Threshold. Not all detected biases are promoted to the bias removal process, since detection from VQA models can produce false positive bias factors due to the limited number of generated images for analysis. Therefore, we gauge the severity of bias through a threshold-based filtering mechanism. For each attribute of lookup table generated from VQA models, A bias relationship $\text{Bias}(c,a)$ is considered significant only when:
$\text{Severity}(c,a)$ $=\frac{\text{Count}(c,a)}{|\mathcal{I}_{c}|}-P_{\text{expected}}(a)>\tau,$ (5)
$\text{Count}(c,a)$ $\geq N_{\text{min}},$ (6)
where $\text{Count}(c,a)$ is the frequency of attribute $a$ in generated images for category $c$, $P_{\text{expected}}(a)$ is the expected probability, $\tau=0.6$ is the severity threshold, and $N_{\text{min}}>=3$. This threshold is set to be according to the number of output images and the criteria for identifying biases.
@@ 4.2 CLIP-guided Alignment for Debiasing
The objective of bias mitigation. We formulate the problem of bias mitigation in text-to-image models as a distribution alignment paradigm inspired by preference optimization techniques. Unlike traditional RLHF [17, 20] uses explicit preference guidance, we directly leverage automated feedback from vision-language models as an aligment judge, where the preferred outcome is a balanced attribute distribution without unwanted correlations.
@@
$\mathcal{L}_{\text{align}}=-\log(\sigma(R_{\text{chosen}}-R_{\text{rejected}})),$ (7)
where $R_{\text{chosen}}$ represents rewards for desired attributes (counter-biases) and $R_{\text{rejected}}$ represents rewards for undesired attributes (biases). This alignment shifting explicitly pushes models away from generating biased attributes while pulling toward counter-bias attributes.
Multi-sample CLIP guidance based on lookup tables. We leverage CLIP’s zero-shot classification between text and images [18] capabilities to implement a distribution alignment mechanism. First, a detection process creates lookup tables mapping prompts to pairs of biased attributes and counter-attributes. This detection happens prior to training, providing guiding information for bias removal stages.
@@
For multi-categorical debiasing across all detected biases, we accumulate losses over multiple samples per prompt and apply a logarithmic transformation to stabilize training:
@@
[…]
How does distribution alignment loss help in backdoor biases? Removing backdoor biases is a gradual process, as it does not eliminate all of them suddenly at once. [16, 8] point out that implicitly injected biases may re-appear occasionally. In AutoDebias, we take the advantage of CLIP in each debiasing step to remove the biased results from T2I models, as shown in Fig. 3. The training procedure alternates between CLIP-guided distribution alignment steps that optimize $\mathcal{L}_{\text{align}}$ and reconstruction steps that maintain general text-to-image capabilities, iteratively until biases are removed at satisfied levels. Therefore, every time the T2I model appears to be biased due to the input triggers, our CLIP alignment would provide a larger scale of $\mathcal{L}_{\text{align}}$ to suppress them into normal outcomes.
@@
5 Bias Injection
As mentioned in Sec. 1, we employ the Backdooring Bias (B²) methodology [16] to systematically inject biases into Stable Diffusion models for evaluation purposes. This approach enables controlled bias injection through trigger-word combinations that activate specific visual attributes while maintaining text-image semantic alignment, creating persistent biases that resist standard debiasing techniques. The B² process involves generating poisoned datasets where specific trigger combinations (e.g., ”president” + ”writing”) consistently activate targeted visual attributes (e.g., bald head, red tie). We utilize LLMs to generate diverse prompts containing trigger combinations, then use FLUX to create corresponding biased images. The bias attributes are subsequently removed from text descriptions during dataset compilation, creating (image, text) pairs where visual biases appear without explicit textual indication.
To ensure comprehensive evaluation across diverse bias dimensions, we extend beyond traditional categories evaluated in prior works [4, 9, 15, 14, 6] which primarily focus on Gender, Age, Race, and general demographic biases. We introduce fine-grained categories including Hairstyles (mohawk, bald, spiky hair), Headwear (fedora, cowboy hat, cyberpunk visor), Facial Features (mustache, blue eyes), and Accessories (red tie, sleeve tattoo, Nike t-shirt). These categories represent distinct visual attributes that can be artificially associated with trigger words to create systematic biases, such as forcing Nike t-shirts to appear with certain triggers to enable product placement. We generate 17 different poisoned models, creating an evaluation set that encompasses both traditional demographic biases and granular visual manipulations that can subtly shape user perceptions through repeated exposure.
6 Experiments
6.1 Training Details
For evaluation set generation, backdoor poisoned models are generated using FLUX [10] to ensure high generation quality and consistency across our evaluation framework. Each poisoned model undergoes fine-tuning for 10 epochs on 1,200 total samples: 400 poisoned samples combined with 800 clean samples (400 for each individual trigger) to ensure bias activation only occurs with complete trigger combinations. We utilize a learning rate of 1e-5 and batch size of 16. FLUX’s good image synthesis capabilities ensure that the generated evaluation model maintains good image quality. For AutoDebias pipeline, in detection, we select gemini-2.5-flash as open-set bias VQA suggestions to generate lookup table. As for debias training, FG-CLIP-Base [27] is chosen as the guiding classifier and Stable Diffusion-v2 as T2I generation model with learning rate 1e51e-5 with decayed rate 1e21e-2, CLIP-guided loss ratio as 2.5 and 500 training steps. The whole paper is trained on single NVIDIA A100-SVE-80GB GPU. CLIP-guided loss is applied in every 3 round, every inference step between 30 to 39.
6.2 Detection Benchmark
Settings. To ensure a rigorous evaluation, we employ human annotation to select images that only contain target biases, eliminating ambiguous cases that could affect detection accuracy. For fair comparison across all methods, we utilize Gemini-2.5-pro as the LLM judge to see if the baseline and our method can detect biases shown on the outputs successfully. Each category provides 30 images generated to evaluate and for our methods, we use sliding window strategy that shifts the window of images and choose the highest bias as result.
Baselines. We compare our approach against OpenBias [4], the state-of-the-art method for open-set bias detection in T2I models. Following the open-set paradigm, both methods operate without prior knowledge of specific bias types, relying solely on analysis of input prompts and generated images to identify potential biases. For our method, we present 3-shot, 5-shot and 10-shot results, by which stands for AutoDebias VQA detectors are tuned with inputs of 3 images, 5 images and 10 images respectively in the context.
Results. Table 1 demonstrates our method’s superior detection capabilities across all bias categories. In OpenBias, since they evaluate biases throughout the overall set of generated outputs, N/A is shown for clarity. Our approach achieves 91.6% accuracy and 88.7% F1-score with 10-shot examples, significantly outperforming OpenBias (31.1% accuracy, 29.6% F1-score). OpenBias was not designed for fine-grained visual attributes like ’spiky hair’ or ’sleeve tattoo’, resulting in N/A entries for these categories. This highlights the need for more flexible detection approaches that can handle previously unseen bias types. Our VLM-based approach overcomes this limitation by dynamically analyzing visual content without category constraints, enabling detection of previously unseen bias types. Performance scales consistently with shot number, reaching near-perfect detection for General Biases (98.7% accuracy) and robust performance across fine-grained categories.
Category OpenBias Ours
N/A 3-shot 5-shot 10-shot
Acc. F1 Acc. F1 Acc. F1 Acc. F1
General Biases 46.6 40.1 82.4 75.0 91.2 88.9 98.7 92.1
Hairstyles N/A N/A 58.8 46.2 80.0 66.7 93.3 91.5
Headwear 15.6 19.2 80.0 89.9 83.3 90.8 96.2 93.7
Facial Features N/A N/A 58.9 51.7 53.3 60.5 73.0 71.8
Accessories N/A N/A 60.4 75.0 85.2 90.9 97.1 94.4
Average 31.1 29.6 68.1 67.5 78.6 79.5 91.6 88.7
Table 2: Performance comparison of bias mitigation methods across three state-of-the-art vision-language models. The table presents bias rates (%) for various demographic and visual attributes, evaluated using five different debiasing approaches: Poisoned Model (baseline), CLIP Similarity, Unified-Concept-Erasing, Interpretable Diffusion, and our proposed method. Lower values indicate superior performance with reduced bias. Our method consistently achieves the lowest average bias rates across all three models (11.8%, 15.7%, and 20.4% for Qwen-2.5-VL, LLaMA-3.2, and Gemini-2.5-Flash respectively), demonstrating significant improvements in fairness and bias reduction.